The following totals are used to create a CVP Income Statement for Frederick Company for FY2018:

Frederick Company

Selected Financial Figures

For the Year Ended 12/31/18

Sales (100 units)

$10,000

Variable Costs:

     Direct Labor

$2,000

     Direct Materials

1,650

     Factory Overhead (variable)

2,000

     Selling Expenses (variable)

600

     Administrative Expenses (variable)

500

Fixed Costs:

     Factory Overhead (fixed)

$800

     Selling Expenses (fixed)

1,000

     Administrative Expenses (fixed)

1,000

Frederick Company utilizes a JIT production system and there are no Raw Materials, Work-in-Process or Finished Goods inventories. Use this information to determine the FY 2016 breakeven point in units. Round and enter as a whole number.
Answer #1

fixed cost = Selling Expenses (fixed) +   Administrative Expenses (fixed)+Factory Overhead (fixed)

= 800+1000+1000

= 2800

Variable Costs: =  Direct Labor+Direct Material+Factory Overhead (variable)+ Selling Expenses (variable)+  Administrative Expenses (variable)

= 2000+1650+2000+600+500

= 6750

variable cost per unit = 6750/100  

= 67.5 per unit

sale price per unit = 10000/100

= 100 per unit

Contribution per unit = sale price - variable cost

= 100-67.5

= 32.5 per unit

break even in units =

fixed cost / contribution per unit

break even in units = 2800/32.5

= 86.15 units

= 87 units